OpenStudy (anonymous):

Brandon is on one side of a river that is 50 m wide and wants to reach a point 200 m downstream on the opposite side as quickly as possible by swimming diagonally across the river and then running the rest of the way. Find the minimum amount of time if Brandon can swim at 1.5 m/s and run at 5 m/s. (Round your answer to two decimal places.)

OpenStudy (mertsj):

\[t=\frac{d}{1.5} + \frac{250-d}{5}\]

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

do you have to use a^2 +b2=c^2 to find the value of d?

OpenStudy (anonymous):

sq200-50^2/1.5

OpenStudy (mertsj):

You need to find the minimum so you should find the first derivative, set it equal to 0 and solve

OpenStudy (anonymous):

there is a derivative in there somewhere needed to be taken too lol

OpenStudy (mertsj):

You need to find the minimum so you should find the first derivative, set it equal to 0 and solve

OpenStudy (anonymous):

so the derivative of \[((\sqrt{200^{2}-50^{2}})/1.5)+250-\sqrt{200^{2}-50^{2}}/5\]